OpenStudy (anonymous):

Evaluate the summation of negative 5 n minus 1, from n equals 3 to 12..

OpenStudy (anonymous):

\[\sum_{n=3}^{12}-5n-1\]

OpenStudy (lxelle):

Integrate the function -5x-1

OpenStudy (anonymous):

How do you do that?

OpenStudy (lxelle):

Go and check out the basic rules of integration. You need to know them well before you do this question.

OpenStudy (lxelle):

\[\int\limits_{?}^{?}x ^{n} dx = \frac{ x^n+1 }{ n+1 }\]

OpenStudy (zpupster):

if you have not studied integration you can do it like this -5(3)-1 = -5(4)-1 = -5(5)-1 = -5(6)-1 = -5(7)-1 = -5(8)-1 = -5(9)-1 = -5(10)-1 = -5(11)-1 = -5(12)-1 = then add up all of these--- integration is faster and easier